How to factor cubic linear equations?
X3-8?

Answer by xcentaur(357)

{X 3 -Y 3 = (X - Y)(X 2 + XY + Y 2 )}


x^3 - 8 = (x - 2)(x^2 + 2x + 4)


x^3 - 8 = (x - 2)[(x-2)^2 + 4x]


x^3 - 8 = (x - 2)^3 + 4x


Hope this helps